回路工学特論 2 単位 (選択)

Advanced Circuit Theory  西尾芳文 · 教授/ システム創生工学専攻電気電子創生工学コース知能電子回路講座

上手洋子 · 助教/ システム創生工学専攻電気電子創生工学コース知能電子回路講座

【授業目的】非線形回路の解析手法，および，非線形振動回路の振る舞いとその

応用について講述する．【授業概要】非線形電子回路の素子のモデリング手法，動作点解析，過渡解析手

法，定常振動解の求解法などについて講述する．次に，非線形振動回路にみられる同期現象やカオス現象などの振る舞いについて講述し，これら非線形振動現象の情報通信工学などへの応用について講述する． (授業形態:講義形式)

【授業形式】講義および演習【キーワード】非線形回路，回路解析，振動回路【先行科目】『電気回路 1· 演習』 (0.7)，『電気回路 2· 演習』 (0.5)，『回路網解析』

(0.3)

【関連科目】『複雑系システム工学特論』 (0.5) 【履修要件】基本的な線形回路の解析手法を習得していること．【履修上の注意】講義は，原則として英語で行う．【到達目標】

1. 非線形回路の解析手法を理解する

2. 非線形回路の振る舞いとその応用について理解する【授業計画】

1. 非線形素子のモデリング

2. Newton 法による非線形回路の動作点解析

3. 非線形回路の過渡解析手法

4. 非線形回路の定常振動解の求解法 (4 回分)

5. 非線形振動回路の同期現象 (2 回分)

6. 非線形振動回路のカオス現象 (2 回分)

7. 非線形回路の工学的応用 (4 回分)

8. まとめとテスト【成績評価基準】試験 80%，平常点 (レポート)20%で評価し，全体で 60%以上あれ

ば合格とする．【教科書】なし【参考書】

⋄ 牛田明夫，田中衞著，「電子回路シミュレーション」，コロナ社 ⋄ スメール，ハーシュ著，「力学系入門」，岩波書店【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216580 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 西尾 (E 棟 3 階南 D-7, 088-656-7470, nishio@ee.tokushima-u.ac.jp) ���� 【備考】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

Advanced Circuit Theory 2 units (selection)

Yoshifumi Nishio · Professor / Intelligent Networks and Computer Science, Electrical and Electronic Engineering, Systems Innovation Engineering Yoko Uwate · Assistant Professor / Intelligent Networks and Computer Science, Electrical and Electronic Engineering, Systems Innovation Engineering

Target⟩ Analysis method of nonlinear circuits, and behavior of nonlinear oscillatory

circuits and its applications are lectured. Outline⟩ Modeling of nonlinear devices, DC analysis and transient analysis of

nonlinear circuits, and steady-state analysis are lectured. Synchronization and chaotic phenomena in nonlinear oscillatory circuits are introduced and their applications to information and communication engineering are discussed. (Style: Lecture)

Style⟩ Lecture and excercise �������⟩ nonlinear circuits, circuit analysis, oscillatory circuits Fundamental Lecture⟩ “Electrical Circuit Theory (I) and Exercise”(0.7), “Elec-

trical Circuit Theory (II) and Exercise”(0.5), “Network Analysis”(0.3) Relational Lecture⟩ “Advanced Theory of Complex System Engineering”(0.5) Requirement⟩ Students are required to know how to analyze basic linear circuits. Notice⟩ Course is taught in English.

Goal⟩

1. Understanding of analysis methods of nonlinear circuits.

2. Understanding of behavior of nonlinear circuits and its applications.

Schedule⟩

1. Modeling of nonlinear devices.

2. DC analysis of nonlinear circuits by Newton method.

3. Transient analysis of nonlinear circuits.

4. Steady-state analysis of nonlinear circuits (for 4 lectures).

5. Synchronization phenomena in nonlinear oscillatory circuits (for 2 lectures).

6. Chaotic phenomena in nonlinear oscillatory circuits (for 2 lectures).

7. Engineering applications of nonlinear circuits (for 4 lectures).

8. Conclusions and final examination. Evaluation Criteria⟩ Final examination 80% and exercise 20%. Textbook⟩ None.

Reference⟩

⋄ Akio Ushida and Mamoru Tanaka, ”Nonlinear Circuit Simulations,” Corona Publishing Co.

⋄ S. Smale and M.W. Hirsch, ”Differential Equations, Dynamical Systems, and Linear Algebra,” Academic Press.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216580

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Nishio (D-7, E-3F-South, +81-88-656-7470, nishio@ee.tokushima-u.ac.jp)

���� Note⟩ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．